A meal consisting of a cheeseburger large fries and chocolate shake provides a total of 1120 kcal. Forty-eight percent of the energy in the meal is from carbohydrate and 13% is from protein. How many calories of fat does the meal contain?
Answer #1

Total amount of calories =1120 Kcal.

Calories in the foods are provided by carbohydrates , proteins and fats.

Given,

Percent of calories provided by carbohydrates = 48

Percent of calories provided by protein = 13

Hence, percent of calories provided by fats = 100 - ( 48+13)

= 39

Therefore amount of calories provided by fat

= ( percentage of fat×total calories in the food) ÷ 100

= (39×1120)÷100= 436.8 Kcal = 436800 cal

Hence, the meal contains 436800 calories of fat.